OpenStudy (anonymous):

hi my name is diamond, I NEED HELP!!!!! HHHHHHHHHHHHHHHHHHHEEEEEEEEEEEEEEEEEEEEELLLLLLLLLLLLLLLLLLLLLLPPPPPPPPPPPPPPPPPP

OpenStudy (yuki):

yes?

OpenStudy (anonymous):

what with :)?

OpenStudy (anonymous):

its adding A LOT of fractions

OpenStudy (yuki):

adding fractions are always the same, let's cope with them one by one

OpenStudy (anonymous):

1/6+2/12+1/3+3/4+1/2=???

OpenStudy (yuki):

the lowest common denominator is = ? if you can answer this then you are almost done

OpenStudy (anonymous):

uuummm 12?

OpenStudy (yuki):

good

OpenStudy (anonymous):

yay

OpenStudy (yuki):

all you have to do is to change all of those fractions so that the bottom is 12 but equivalent, for example 1/6 = 2/12

OpenStudy (anonymous):

add the numorators? thats how my teach tought me

OpenStudy (yuki):

before you add them, all fractions have to have the denominators the same

OpenStudy (anonymous):

ohohoh, 23/9???

OpenStudy (yuki):

so once your fractions have all the same denominator,you can start adding your numerators

OpenStudy (anonymous):

um 23/9

OpenStudy (pleasehelpme):

yes

OpenStudy (yuki):

the denominator is 12, remember?

OpenStudy (anonymous):

yay!!!

OpenStudy (anonymous):

oh yeah 23/12

OpenStudy (yuki):

that's the right one emo, never fear fractions from now on :)

OpenStudy (anonymous):

kk thank u